OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (0, 7) and (5, 3)

OpenStudy (aroub):

Slope: 3-7/5-0 -4/5 Equation: y=mx+b , m=-4/5 and ( 0,7) 7=-4/5(0)+b 7=b => y=-4/5x+7

OpenStudy (anonymous):

thannk you soo muchh :))

OpenStudy (anonymous):

y=-(4/5)x+7